OpenStudy (anonymous):

Evaluate the given integral using the substitution (or method) indicated. (x + 1)e^(x + 1)^2 dx; u = (x + 1)2

jimthompson5910 (jim_thompson5910):

if u = (x+1)^2, then du/dx = 2(x+1) which means du = 2(x+1)*dx the original integrand (x + 1)e^(x + 1)^2 dx has (x+1) and dx in it, so isolate (x+1)dx in du = 2(x+1)*dx to get du = 2(x+1)*dx du/2 = (x+1)*dx (x+1)*dx = du/2 with me so far?

jimthompson5910 (jim_thompson5910):

or no?